how do i find two consecutive odd positive integers with a product of 399??
Answer by JimboP1977(308)

let x be a odd integer
let x+2 be the next odd integer
x * (x+2) = 399
x^2 +2x = 399
(x+1)^2 - 1^2 = 399
x = sqrt(399+1)-1
x = 20-1
x = 19
x+2 = 21
Proof
21*19=